OpenStudy (anonymous):

how would you turn 9x=177+4y into x+y=

OpenStudy (anonymous):

someone please help me

OpenStudy (anonymous):

thats not what im looking for.

OpenStudy (anonymous):

if its 9x=177+4y it needs to look something like 9x+4y=177

OpenStudy (anonymous):

9x+4y=177 -4+177=173 9x=173y 173y/9=19.2y x=19.2y -19.2x=19.2y -19.2y+19.2x=x+y there you have it...